1 3 5 7 9 11 13 15 17 19, To Make 50 Use Only 5 Digits ...? What The 5 Digits
Found 2 solutions by Fombitz, ikleyn:
Answer by Fombitz(32388)

If you mean by addition/subtraction, it can't be done.
Two odds numbers sum to an even number.
So two sets of odd numbers would sum to an even number.
That's used up 4 of the 5.
Adding another odd number makes the sum odd.
Example:
1+3=4, Even
5+7=12, Even
4+12=16, Even
So then
16+19=35, Odd

Answer by ikleyn(52779)

.
How the post is presented and worded, it makes no sense.